OpenStudy (anonymous):

how t solve C/4 - 3 =8

OpenStudy (anonymous):

C/4 - 3 =8 C/4 - 3+3 =8+3 C/4 = 11 C=44

OpenStudy (anonymous):

so isolate 3 then multiply or divide 4 & 11?

OpenStudy (anonymous):

As we want to find C, the final answer is expressed into C=..... so make C as the subject. we want to cancel -3 on LHS so +3 to both sides then we want to cancel 4 on LHS so multiply 4 to both sides

OpenStudy (anonymous):

ok cool, thanks for explanation. I am practicing for exams so I need all the help i can get

OpenStudy (anonymous):

What a coincidence! I am practicing for my exams too :D